How much Oil (in bbl), Water (in bb) and Hematite (in sacks) are required to make exactly 1000 bbl of 16 ppg oil mud with an oil water ration of 90/10. Assume an oil density of 6.8 ppg and a Hematite density of 42 ppg
Answer #1

Volume of oil =V1

Volume of water=V2

Volume of Hematite =V3

Density

w1=6.8 ppg

water density in ppg=8.33 ppg

so

w2=8.33 ppg

w3=42 ppg

So for adding theses three component we find new oil mud.

So mass =density*volume

So v1w1+v2w2+v3w3=vfwf

The oil and water ratio is 90/10.So that they can be considered as one volume v that is 90% of oil and 10% water.

v1=0.9v

v2=0.1v

v1+v2+v3=1000

v+v3=1000

So

v3=1000-v

So put all values in v1w1+v2w2+v3w3=vfwf.

0.9*v*6.8+0.1*v*8.33+(1000-v)*42=1000*16

35.047v=26000

v=741.86 bbl

so oil =0.9*v=667.67 bbl

water=0.1*v=74.186 bbl

hematite=1000-v=258.14 bbl